# A search for transit timing variations in the transiting hot Jupiter systems HIP 65, NGTS-6, NGTS-10 and WASP-173

### v1

@@ Abstract
Hot Jupiters are Jupiter-mass planets with orbital periods of less than ten days. Their short orbital separations make tidal dissipation within the stellar host especially efficient, potentially leading to a measurable evolution of the orbit. One possible manifestation of this is orbital decay, which presents itself observationally through variations in the orbital period and thus times of transit. Here we select four promising exoplanetary systems for detecting this effect: HIP 65, NGTS-6, NGTS-10 and WASP-173. We present 33 new transit light curves taken with the 1.54 m Danish Telescope, and analyse these alongside photometric data from the Transiting Exoplanet Survey Satellite and transit timing data from the literature. We construct two ephemeris models for each target: a linear ephemeris and a shrinking orbital period due to tidal decay. The linear ephemeris is preferred for all four models – the highest significance for the quadratic ephemeris is 2 for WASP-173. We compare these results to theoretical predictions for tidal dissipation of gravity waves in radiation zones, and find that wave breaking is predicted only in WASP-173, making rapid decay plausible in this system but unclear in the other three. The sensitivity of transit timings to orbital decay depends on the square of the time interval covered by available observations, so our results establish a useful baseline against which future measurements can be compared. NGTS-6 and NGTS-10 are important objects for future study as they are in the first field to be observed by the upcoming PLATO mission.
@@ 4.1 HIP 65 Ab
From the orbital decay model, we find a positive quadratic coefficient, corresponding to a period increase of . For comparison, the published period derivatives for HIP 65 Ab are from 2024PSJ.....5..163A and a rate of period change per cycle of by 2024A&A...692A..35M. Our result is consistent with these but also with a constant period. The statistical analysis shows no substantial favouring of either model, with BIC = -1.4 and AIC = 0.0, acting as sufficient reason not to explore models with higher-order polynomials. The detection of any orbital decay in this system would require observations over a longer time span than currently available.
@@ 4.2 NGTS-6 b
We find the difference between the strength of the linear and quadratic models to be BIC = 3.6 and AIC = 4.9, which does not significantly favour the quadratic model. Alongside an inspiral rate of , tidal decay is not significantly favoured over the constant-period alternative. Although the best-fitting value is negative, this result is consistent with zero at 1.2 and therefore does not directly imply a shrinking orbit. Further observations will be paramount in distinguishing orbital decay from typical orbital motion. Opportunely, NGTS-6 b is in a sample of known transiting exoplanets that will be observed by the PLATO mission (Rauer+25exa) in its first long-pointing field (LOPS2; 2025A&A...694A.313N). As a result, the occasion may present itself in less than a decade with the release of PLATO data. This planet makes an interesting tidal decay candidate for future study.
@@ 4.3 NGTS-10 b
A successful transit timing analysis of NGTS-10 is hindered by the lack of available transit times. Although the times used in this study span over eight years, only 12 transit times have been obtained. The value presented here is negative with a significance of just above 1 based on the available data. There is a gap of several years between the earlier times and the observations presented in the current work (see Fig. 2), meaning that a linear ephemeris is precisely anchored but that the quadratic term is not well constrained.
@@ 4.4 WASP-173 Ab
Our timing analysis of WASP-173 Ab produced a lower and than those detected in other works, with 2024PSJ.....5..163A finding and 2024A&A...692A..35M finding . The differences are likely attributed to the variety of data sources used in each work. We find BIC = 12.3 and AIC = 13.2, indicating the quadratic model fits the residuals better than that of the linear. is only negative at 1.9, so the orbital decay scenario is still not conclusive. However, 2024A&A...692A..35M predict wave breaking provided the stellar host is older than , giving the theoretical constraint which agrees with our own theoretical estimation. As with the other three systems, more timing data spanning a longer time interval will be needed to identify orbital decay in WASP-173 A.
@@ 5 Conclusions
By taking the 3 lower limits of for each system, we computed lower limits for . These are not remarkable, lying in the range , but are compatible with our theoretical predictions if tidally-driven orbital decay due to gravity waves is actually occurring. Given that the current constraints are not very far from predictions, each of these 4 systems would be promising for future follow-up studies since they have very good potential to test tidal theory. On the other hand, the possible outward migration of HIP 65 Ab, if confirmed in future studies, would be more difficult to explain by tidal processes alone in such a slowly rotating star. A significant positive , if found, would be most likely explained by another phenomenon, such as the light time effect (e.g. 2025ApJ...986..117Y) or an acceleration of the system away from the observer. As of now, our TTV models show no conclusive evidence for orbital decay in any of these systems. WASP-173 Ab has the strongest suggestion of orbital decay and is a promising candidate for future studies.

### v2

@@ Abstract
Hot Jupiters are Jupiter-mass planets with orbital periods of less than ten days. Their short orbital separations make tidal dissipation within the stellar host especially efficient, potentially leading to a measurable evolution of the orbit. One possible manifestation of this is orbital decay, which presents itself observationally through variations in the orbital period and thus times of transit. Here we select four promising exoplanetary systems for detecting this effect: HIP 65, NGTS-6, NGTS-10 and WASP-173. We present 33 new transit light curves taken with the 1.54 m Danish Telescope, and analyse these alongside photometric data from the Transiting Exoplanet Survey Satellite and transit timing data from the literature. We construct two ephemeris models for each target: a linear ephemeris and a shrinking orbital period due to tidal decay. The linear ephemeris is preferred for three of the four models – the highest significance for the quadratic ephemeris is over 3 for WASP-173. We compare these results to theoretical predictions for tidal dissipation of gravity waves in radiation zones, and find that wave breaking is predicted only in WASP-173, making rapid decay plausible in this system but unclear in the other three. The sensitivity of transit timings to orbital decay depends on the square of the time interval covered by available observations, so our results establish a useful baseline against which future measurements can be compared. NGTS-6 and NGTS-10 are important objects for future study as they are in the first field to be observed by the upcoming PLATO mission.
@@ 4.1 HIP 65 Ab
From the orbital decay model, we find a positive quadratic coefficient, corresponding to a period increase of . For comparison, the published period derivatives for HIP 65 Ab are from 2024PSJ.....5..163A and a rate of period change per cycle of by 2024A&A...692A..35M. Our result is consistent with these but also with a constant period at under 2. The statistical analysis shows no substantial favouring of either model, with and , acting as sufficient reason not to explore models with higher-order polynomials. The detection of any orbital decay in this system would require observations over a longer time span than currently available.
@@ 4.2 NGTS-6 b
We find the difference between the strength of the linear and quadratic models to be and , which does not significantly favour the quadratic model. Alongside an inspiral rate of , tidal decay is not significantly favoured over the constant-period alternative. Although the best-fitting value is negative, this result is consistent with zero at 2.3 and therefore does not directly imply a shrinking orbit. Further observations will be paramount in distinguishing orbital decay from typical orbital motion. Opportunely, NGTS-6 b is in a sample of known transiting exoplanets that will be observed by the PLATO mission (Rauer+25exa) in its first long-pointing field (LOPS2; 2025A&A...694A.313N). As a result, the occasion may present itself in less than a decade with the release of PLATO data. This planet makes an interesting tidal decay candidate for future study.
@@ 4.3 NGTS-10 b
A successful transit timing analysis of NGTS-10 is hindered by the lack of available transit times. Although the times used in this study span over eight years, only 12 transit times have been obtained. The value presented here is negative with a significance of just above 2 based on the available data. There is a gap of several years between the earlier times and the observations presented in the current work (see Fig. 2), meaning that a linear ephemeris is precisely anchored but that the quadratic term is not well constrained.
@@ 4.4 WASP-173 Ab
Our timing analysis of WASP-173 Ab produced a lower and than those detected in other works, with 2024PSJ.....5..163A finding and 2024A&A...692A..35M finding . The differences are likely attributed to the variety of data sources used in each work. We find and , indicating the quadratic model fits the residuals better than that of the linear. Although is negative at 3.8, this result could be optimistic, caused by underestimated uncertainties on some of the transit times. The orbital decay scenario is still not conclusive. Nevertheless, 2024A&A...692A..35M predict wave breaking provided the stellar host is older than , giving the theoretical constraint which agrees with our own theoretical estimation. As with the other three systems, more timing data spanning a longer time interval will be needed to identify orbital decay in WASP-173 A.
@@ 5 Conclusions
By taking the 3 lower limits of for each system, we computed lower limits for . These are not remarkable, approximately lying in the range , but are compatible with our theoretical predictions if tidally-driven orbital decay due to gravity waves is actually occurring. Given that the current constraints are not very far from predictions, each of these 4 systems would be promising for future follow-up studies since they have very good potential to test tidal theory. On the other hand, the possible outward migration of HIP 65 Ab, if confirmed in future studies, would be more difficult to explain by tidal processes alone in such a slowly rotating star. A significant positive , if found, would be most likely explained by another phenomenon, such as the light time effect (e.g. 2025ApJ...986..117Y) or an acceleration of the system away from the observer. As of now, our TTV models show no conclusive evidence for orbital decay in any of these systems. WASP-173 Ab has the strongest suggestion of orbital decay and is a promising candidate for future studies.